### v1

# Dynamical Schwinger Effect in Lead-Halide Perovskites

Abstract
We observe strong photoluminescence of a lead-halide perovskite driven by deep sub-gap irradiation. Using the quasi-adiabatic Landau-Dykhne approach, we interpret this observation in terms of the dynamical Schwinger effect – tunneling ionization in strong fields. Further, the exponential sensitivity of the Schwinger effect to driving fields allows us to measure the local frozen-in fields in a nominally cubic MAPbBr3 single crystal at room temperature. Finally, we demonstrate an AC analogue of biasing in our system – the cooperation between two time-dependent fields simultaneously driving the sample. Our results (i) establish lead-halide perovskites as an excellent platform for simulating effects of strong fields on Dirac fields, (ii) contribute to the on-going discussion about inversion-breaking in MAPbBr3 single crystal, and (iii) pave the way for a mid-infrared light detection with lead-halide perovskites.
I Introduction
One of the most important insights coming from the synthesis of Quantum Mechanics and Special Relativity was the realization that vacuum is not empty. A textbook demonstration of this remarkable fact would be an observation of an electric breakdown of an empty space. However, as was shown by Schwinger, this can only occur at extremely strong fields $E_{S}\approx 10^{20}$/m Schwinger (1951), which are currently beyond experimental reach. In view of the enormity of $E_{S}$ and its intuitive interpretation as the maximal static electric field attainable in principle, it is natural that the actual attempts to realize vacuum breakdown (“Schwinger Effect”, SE) focus on its dynamical versions where the vacuum pair creation is achieved under the influence of time-dependent fields Mourou et al. (2006). The experimental progress is motivated by a number of theoretical works that suggest to reduce the field amplitudes needed for the observation of SE by crafting profiles of the driving fields Popov (1971); Ringwald (2001); Schützhold et al. (2008). However, any theoretical proposal must rely on certain simplifying assumptions, which due to the inherently non-perturbative character of SE may be decisive to the outcome Gies and Klingmüller (2005); Torgrimsson et al. (2017). Given that current technology is not capable to test these ideas in practice, it is hard to tell at the moment how realistic will they ultimately turn out to be.
In this context it appears desirable to be able to simulate the Schwinger effect and the various approaches to it in a more accessible setting. An obvious direction is a realization of field-induced tunneling ionization in a semiconductor with a Dirac-like band dispersion. It is straightforward to demonstrate then that the threshold field in this case would be $E_{s}\sim\Delta/ea\sim 10^{9}$/m Linder et al. (2018), which is readily achievable in modern laboratories (here $\Delta$, $a$ and $e$ are the band gap, unit lattice length and electron charge respectively). However, it turns out that the observation of this scaled-down version of SE also poses significant challenges. The main difficulty here lies in the fact that despite an abundance of far-reaching analogies Wehling et al. (2014), there is an important difference between the “true” Dirac field in vacuum and a Dirac material in that the former features relativistic invariance while the latter does not. The important consequence in question is that a free particle in vacuum can in principle be accelerated indefinitely, while a charge carrier in a semiconductor can produce secondary particles once it acquires a certain threshold kinetic energy comparable to the band-gap Jena (2022). Since the threshold for this process $E_{\mathrm{av}}\sim\Delta/e\lambda_{\mathrm{mfp}}$ is significantly lower than that for tunneling ionization (here $\lambda_{\mathrm{mfp}}>a$ is the mean free path of charged carriers in the material), the population of secondary charge carriers will tend to dominate over tunnel-ionized carriers, often ending up in a catastrophic avalanche breakdown of the material. As a result of these complications, the observations of tunneling ionization have been limited so far either to the cases where the relevant tunneling region is restricted to a narrow part of space to avoid avalanche breakdown (e.g. tunneling ionization in single atoms Raizer (1991) or tunnel diodes Esaki (1958) and tunnel junctions Giaever and Megerle (1961); Chen (2007) in electronics); or to extended systems with no gap Berdyugin et al. (2022). However, to establish a direct connection to the efforts in high energy physics, it is necessary to realize an analog Schwinger effect in an extended Dirac material with a finite bandgap.
In this Article, we report the observation of dynamical tunneling ionization in a single-crystal sample of lead halide perovskite (LHP), which realizes a gapped Dirac system Jin et al. (2012). In order to suppress avalanche formation and emphasize tunneling, we note that the former needs a finite amount of time to build up while the latter occurs quasi-instantaneously. Therefore, we seek to induce ionization with alternating electric fields using frequencies that, on the one hand, correspond to photon energies significantly smaller than the band gap $\hbar\omega\ll\Delta$ (to ensure that we operate in the quasi-adiabatic regime); on the other hand, they should oscillate fast enough to avoid an avalanche breakdown ($\omega>eE/\sqrt{m\Delta}$, where $E$ and $m$ are the magnitude of the applied electric field and the band mass of the charge carrier, respectively). Additionally, we are careful to keep our irradiation intensity levels well below impact ionization thresholds for our pulse durations $\tau\approx 300$fs to suppress the population of secondary charge carriers Rethfeld et al. (2017). This is necessary to reveal the intrinsic tunneling processes in the material and also to be able to describe them using quasi-adiabatic methods.
Figure 1: A) Schematic diagram of ionization across the energy gap $\Delta$ due to multiphoton process, and due to tunneling under the influence of different applied field amplitudes ($E_{1}$ and $E_{1}+E_{2}$). It can be seen that the magnitude of $E$ determines the width of the forbidden range, thus affecting the net tunneling rate exponentially (see the text for details); B) Tunneling under the influence of the of alternating fields $E_{1}(t)$ and $E_{2}(t)$ (yellow and red); due to exponential sensitivity of the tunneling rate (blue) on the magnitude of the field, the tunneling rate peaks at the maximum of the total field (magenta); consequently the total number of transitions (green) depends only on the sum of the amplitudes of applied time-dependent fields; C) PL spectrum of MAPbBr3 single crystal sample pumped by $\lambda=4\mu$m radiation; Inset: MAPbBr3 crystal with PL coming from the bulk of the sample; D) PL spectra under $\lambda=4\mu$m pumping, together with our theoretical prediction for weak fields (no fit parameters), and for intermediate electric fields (see Supplementary Material). Note that we normalize all curves to be unity at a given value of the field ($eaE_{AC}\simeq 0.272$eV), for convenience. E) An illustration of several most common channels of photo-carrier recombination with characteristic rate dependencies on the total number of charge carriers (bottom).
Low irradiation intensities, however, imply that all measurements need to be performed in the regime where the concentration of tunnel-induced carriers is so low that standard transport measurements cannot be used for studying tunnel ionization. Therefore, our approach here is to infer the population of photo-excited carriers from the intensity of photoluminescence (PL) emitted upon their recombination with each other. The success of this approach crucially depends on the photoluminescent quantum efficiency of the parent material, namely the ratio of radiative recombinations to the total total number of excitations Braslavsky (2007). This consideration is the primary motivation for the choice of lead halide perovskites as the base for realizing the solid-state version of SE. Furthermore, effective behavior of lead halide perovskites renowned for excellent quantum optoelectronic properties Goetz et al. (2020) is described very accurately by a Dirac equation with a non-zero mass Becker et al. (2018); Volosniev et al. (2023a, b), strengthening the analogy between our findings and the physics of SE.
II Observation of Dynamical Schwinger Effect in MAPbBr3
Lead-halide perovskites are known to feature strong sub-bandgap response in terms of two- and three-photon adsorption Zhou et al. (2021); Mei et al. (2022) whose analysis has been mainly focused around the photophysical properties of LHPs in the near-infrared regime as it can shed light onto surprising solar cell (photovoltaic) performance of LHPs Jena et al. (2019). At the same time, the response of the system to photons with lower frequency remains largely unexplored. In Fig. 1C we show the PL spectrum of MAPbBr3 illuminated with mid-infrared radiation ($\lambda\approx 4\mu$m). Given the wavelength of PL ($\lambda_{\mathrm{PL}}\approx 570nm$) which roughly corresponds to the bandgap energy in MAPbBr3 ($\Delta\approx 2.3$eV Saidaminov et al. (2015); Tilchin et al. (2016)) this process corresponds to a whopping frequency conversion factor of $\sim 10$ at the face value. Combined with the exponential sensitivity of this process to the intensity of the irradiating field in Fig. 1D, we are compelled to rationalize the observed PL under deep sub-gap irradiation as quasi-adiabatic tunnel ionization of electrons across the semiconductor band gap using the standard terminology of strong-field phenomena Kruchinin et al. (2018) (similar behavior was observed in CsPbBr3, which indicates the general character of the phenomenon; see Supplementary Material). This observation comprises the central finding of this work. In the following we will further substantiate this qualitative observation by a quantitative analysis of ionization in a periodically driven Dirac-like band structure.
II.1 Quasi-adiabatic tunneling
Tunneling is the quintessential quantum effect whereby a system undergoes a transition that can be described semi-classically by a trajectory of which at least a part passes through a classically-forbidden region in the parameter space. As a result the rate of this transition is strongly suppressed, being exponentially sensitive to the width of the classically-forbidden region Landau and Lifshitz (1981), see Fig. 1A. For instance, in the case of static electric field $E$ the rate of tunneling ionization $W$ across the gap $\Delta$ can be shown to be
$W\propto\exp\left(-E_{s}/|E|\right)$ (1)
with the characteristic cut-off field introduced in the previous sections $E_{s}\sim\Delta/ea$ Sauter (1931); Zener (1934); Schwinger (1951). A convenient way to extend this limiting expression to the problem of tunneling excitation of a semiconductor under the influence of time-dependent fields varying at the characteristic frequencies much smaller than those given by the bandgap is the so-called quasi-adiabatic Landau-Dykhne method Delone and Krainov (2000). Within this approach the ionization rate $W$ can be shown (see Supplementary Material) to be given as
$W\sim\mathrm{exp}\left(-2f(\gamma_{K})\,\Delta/\omega\right),$ (2)
where $f(x)=\mathrm{arsinh}(x)-(\sinh(2\mathrm{arsinh}(x))-2\mathrm{arsinh}(x))/(8x^{2})$, and $\gamma_{K}$ is the so-called Keldysh parameter of the problem whose numeric value depends on the strength and the frequency of the driving field as well as on the details of the band structure of the material in question Keldysh (1965). In the specific case of lead-bromide perovskites, density functional theory calculations provide $\gamma_{K}\simeq 0.85\omega/(eaE_{AC})$, where $E_{AC}$ is the strength of the electric field induced by the laser, and $a\simeq 0.586$nm is the lattice constant MAPbBr3 in cubic phase Becker et al. (2018).
The value of Keldysh parameter determines the degree of adiabaticity of the problem, the two limiting cases being pure adiabatic tunneling ($\gamma_{K}\to 0$) and multiphoton absorption ($\gamma_{K}\to\infty$) as illustrated in Fig.1A. In our experiment $\gamma_{K}\lesssim 1$ (see Fig. 1D), which puts us in the quasi-adiabatic tunneling regime, where the tunneling rate can be well approximated by $W\sim\exp\left[-{1.4\Delta}/{(eaE_{AC})}\right]$ exhibiting the exponential dependence on the external electric field characteristic of the non-perturbative tunneling processes such as (static) Schwinger effect [cf. Eq. (1)].
II.2 Comparison to the experiment
In this section we confirm the tunneling nature of the ionization by deriving the functional dependence of the PL intensity, $I_{\mathrm{PL}}$, on the strength of the pumping mid-infrared (mid-IR) radiation. Naively, one might assume $I_{\mathrm{PL}}\propto W$, i.e., every single tunneling event results in a PL photon. However this is only true if: 1) there are no non-radiative recombination channels; 2) the photo-excited electron-hole pairs remain well-separated in space. Neither of these conditions are satisfied in lead-halide perovskites. First of all, despite the excellent photoelectronic properties of LHPs the non-radiative recombination processes in them cannot be ignored altogether. Similarly, the long diffusion lengths of photocarriers in LHP imply strong overlap between electron-hole pairs which gives rise to a almost exclusively bi-molecular character of the recombination processes Richter et al. (2016); Stranks (2017).
In view of the discussion above, to quantitatively describe the photoluminescence in perovskites, one needs to know the population dynamics of photoexcitations $n(t)$ from which one can calculate the PL intensity based upon the bi-molecular recombination in lead halide perovskites, $I_{PL}\sim\int_{0}^{t}n(t)^{2}\mathrm{d}t$. The general equation governing the population of charge carriers $n(t)$ can be written as:
$-dn(t)/dt=A_{1}n+A_{2}n^{2}+A_{3}n^{3}+....$ (3)
Here, the coefficients $\{A_{i}\}$ describe the rates of different decay channels of $n(t)$ such as: defect-assisted ($A_{1}$), bi-molecular (both radiative- and non-radiative; $A_{2}$), and Auger-type ($A_{3}$) recombination processes followed by higher-order processes which should in principle be taken into account for sufficiently high densities of charge carriers, see Fig. 1E.
In the limit of weak pumping fields, $n(t)$ remains small at all times, implying that $n(t)=n(0)\exp\left(-A_{1}t\right)$ and $I_{\mathrm{PL}}\sim n(0)^{2}$ where the initial population $n(0)$ is proportional to the tunneling ionization rate $W$. As the pumping intensity grows, the omission of higher-order terms in Eq. (3) is no longer justified. For example, the inclusion of both mono- and bi-molecular channels leads to $I_{PL}\sim\alpha W-\mathrm{ln}(1+\alpha W)$, where $\alpha$ is a fitting parameter (see Supplementary Material). This expression in turn is accurate if the densities of charge carriers are of the order of $10^{16}-10^{17}$cm-3 (see Supplementary Material), which occurs for $eaE_{AC}\lesssim 0.3$eV, providing an estimate for a number of excited particles in the system. For higher excitation densities, the channels beyond bi-molecular recombination have to be included.
In Fig. 1D we show the results of the fit limited to second (bi-molecular) processes (yellow curve) as well a phenomenological fit including higher-order channels (red curve) line. The phenomenological fit is motivated by our analytical results. It has the form $I_{PL}\sim\alpha\tilde{W}-\mathrm{ln}(1+\alpha\tilde{W})$, where $\tilde{W}=W/(1+\beta W)$ is a renormalized ionization rate that takes into account that for large densities only a fraction of charge carriers can recombine radiatively. Note that the phenomenological fit has now two parameters $\alpha$ and $\beta$. The overall prefactor that connects $I_{PL}$ to $n^{2}$ is beyond the Landau-Dykhne approach. One could in principle introduce this prefactor as another fit parameter. We refrain from doing so by properly normalizing the data, see Fig. 1D.
III Discussion
[…]
III.1 Tunneling enhancement in LHP due to local frozen-in electric fields.
The distinctive exponential sensitivity of tunneling ionization rate on the peak value of the net applied field can be used as a resource to detect weak local fields in a material. This can be best illustrated in the quasi-static regime describe above. Using the Landau-Dykhne approach it can also be demonstrated in a general case that even a weak static frozen-in field $E_{1}$ ($\omega_{1}=0$) can strongly enhance the tunneling rate $W$ due to the additional driving field $E_{2}(t)$ (see Supplementary Material). The fact that local fields can enhance PL yield is consistent with a qualitative observation that we see stronger PL in the vicinity of structural defects (edges, surfaces) of LHPs where internal electric fields, $E_{1}$, are expected to appear Niesner et al. (2016).
As an application of this sensitivity of PL to $E_{1}$ we will use it to study the highly debated frozen-in electric fields $E_{DC}$ in LHPs Ambrosio et al. (2022), which lead to a broken inversion symmetry even for nominally cubic lead-halide perovskites. It is worth noting here that the method merits certain advantages as compared to some of the conventional techniques for detecting local breaking of inversion symmetry such as second-harmonic generation or angle-resolved photoemission. First of all, our method is not confined to the vicinity of the sample sufrace since the we only use sub-gap frequencies. Second, the peculiarity of tunneling is that it is sensitive to the absolute value of field amplitudes, cf. Eq. (1). Therefore there is a finite net effect due to local inversion breaking even if the local fields spatially average to zero. Finally, another implication of the field cooperation is that PL should strongly depend on the polarization of the external field.
Figure 2: A) PL as a function of the applied electric field for two orthogonal polarization orientations aligned with the crystal axes of MAPbBr3 (markers). The corresponding fits according to our model to the data are presented as solid curves. The inset shows a cartoon of MAPbBr3 with hypothetical ferroelectric domains overlaid with a schematic representation of the irradiating laser spot whose diameter is $d\approx 200\mu$m ($1/e^{2}$); B) Polarization scan taken at a fixed value of the external electric field $eaE_{AC}\simeq 0.180$.
For experimental validation, we measure PL from a single-crystal MAPbBr3 as a function of intensity of the driving mid-infrared $\lambda=4\mu$m radiation. The resulting dependence plotted in Fig. 2A indicates that the tunneling ionization efficiency depends strongly on the polarization of the pumping field. To elucidate the nature of this polarization dependence we scan PL as a function of polarization at a fixed pump intensity as shown in Fig. 2B.
The two-fold symmetric polarization dependence in Fig. 2B might appear surprising since this measurement was performed at room temperature where MAPbBr3 is expected to be in the cubic phase, i.e., four-fold symmetric (e.g., consider the two-photon absorption experiments like in Ref. Li et al. (2020)). The data in Fig. 2B therefore indicates an unexpected lowering of symmetry in MAPbBr3 in the nominally cubic phase. In the most straightforward explanation this could be attributed to an extrinsic factor such as residual stress in the crystal structure that would lift the four-fold symmetry of the electronic band structure. However it can be demonstrated that, that in this case in order to reproduce the magnitude of the effect in Fig. 2 the relative anisotropy of microscopic Hamiltonian describing the system has to reach unrealistic levels of $\sim 50\%$ (see Supplementary Material).
[…]
In an alternative scenario, the observed anisotropy can be attributed to the local frozen-in electric fields, that were conjectured to be present in lead-halide perovskites Niesner et al. (2016); Rakita et al. (2017). Indeed the exponential sensitivity of PL on applied fields implies that one can obtain the anisotropy observed in Fig. 2 already with moderate static fields. In Figs. 2A and B we show that all of the experimental data can be reproduced within the two-field Landau-Dykhne formalism (see Supplementary Material). Based on the fitting, we estimate the numeric value for the internal static fields in room-temperature single-crystal MAPbBr3 to be of the order of $E_{DC}\sim 0.1$V/nm. Within this interpretation of the PL anisotropy, the PL gets maximum enhancement when the polarization of the driving optical field coincides with the direction of $E_{DC}$. That this direction happens to be aligned with direction of Pb-Br bonds ([001] of the cubic lattice) is consistent with previous observations of possible ferroelectricity in MAPbI3 Röhm et al. (2019); Leonhard et al. (2019).
Importantly, we also observe that the two-fold symmetric pattern of PL is subject to the spatial extent of the focal spot of the driving laser. Namely, we observe that that the two-fold symmetry in Fig. 2 is only present for sufficiently focused beams. For larger illumination spots, the pattern turns out to be four-fold symmetric, which can be naturally explained if more than one ferroelectric domain with different field directions participate in the fluorescence process (see Supplementary Material). This interpretation agrees well with the previous estimations of the size of possible ferroelectric domains of several microns in MAPbI3 Rothmann et al. (2017); Rakita et al. (2017).
III.2 Cooperation of dynamic fields and AC-biasing
The hallmark feature of quantum tunneling is its exponential sensitivity to external parameters, which has allowed applications defying the classical common sense. The most celebrated example is the scanning tunneling microprobes where exponential sensitivity of the tunneling current means that the overwhelming majority of it is flowing through the single atom that happens to stick out the most toward the sample  Binnig et al. (1982); Chen (2007). Realization of this has led to unprecedented sub-atom-scale spatial resolutions. Other celebrated applications of tunneling-related process include the tunnel diode Turner (1976) with its unique I-V characteristics and high-harmonic generation in atomic optics where deeply sub-threshold laser light promotes electrons in an atom to unoccupied states Yue and Gaarde (2022), thus initializing the process of extreme frequency conversion of radiation.
[…]
In Fig. 3A we show the cumulative PL coming from the sample as a function of the delay time between the two pulses in cross- and parallel polarization geometries (yellow x- and blue o-markers, respectively). As can be clearly seen, the PL signal is enhanced only when the two pulses of the same polarization overlap in time, i.e., when the amplitude of the sum field is maximized in full agreement with the concept of the dynamical assistance in quantum tunneling introduced at the beginning of this section. As a further illustration of the same phenomenon, we present in Fig. 3(B-D) the 2D dependence of PL on the intensity of each infrared beam.
As a demonstration of the practical use of this effect, we present in Fig. 3E the effect of AC-biasing on PL produced by a mid-infrared beam ($4.5\mu$m) in MAPbBr3333 biased with $1\mu$m radiation at $I_{\mathrm{bias}}=5\times 10^{13}$/m2. Here, we show the differential photoluminescence for two beams $\Delta PL(I)=PL(I_{1},I_{\mathrm{bias}})-PL(I_{1},0)-PL(0,I_{\mathrm{bias}})$, where $PL(I_{1},I_{\mathrm{bias}})$ stands for the total PL emitted by the sample when irradiated by mid- and near-infrared pulses overlapping in time. As can be seen the differential sensitivity of PL can be made linear in the intensity of the mid-IR which opens broad possibilities for efficient mid-infrared sensors based upon lead-halide perovskites. We demonstrate the latter by overlapping the two beams (mid- and near-infrared) on a screen covered by LHP microparticles. As can be seen in Fig. 3F-G, there is clear enhancement in local PL from the region of spatial overlap. Curiously, due to the nonlinear sensitivity of PL on the pumping beam intensity, the apparent size of the overlap spot is significantly smaller than the spot sizes measured for both beams, which can be employed for super-resolution microscopy.
IV Summary
To summarize, we have demonstrated dynamical Schwinger effect in a Dirac semiconductor (lead-halide perovskite) and quantitatively described it with the quasi-adiabatic Landau-Dykhne approach. Further, we have employed the characteristic exponential sensitivity of tunneling ionization on driving fields to measure the local frozen-in fields in the nominally cubic MAPbBr3 single crystal at room temperature contributing to the ongoing debate on the state of ferroelectricity in lead-halide perovskites. Finally, we have investigated the cooperation between two time-dependent fields simultaneously driving tunneling ionization and demonstrated this cooperation can act as an AC analogue of biasing for an optical frequency upconversion, which paves the way for a mid-infrared light detection with lead-halide perovskites.